OpenStudy (anonymous):

(2r-s)/(-4)(r+s)for r=1 and s=-2

OpenStudy (mokeira):

Substitute \[\frac{ ([2\times1]--2 }{(-4)(1-2 )}=\frac{ 2+2 }{-4(-1) }=\frac{ 2+2 }{4 }=simplify\]

OpenStudy (anonymous):

thanks